Using the method of superposition, determine the deflection of the beam shown at point C. - B L a
Answer #1

Question 7 given ay The ic ookie- find of def section at superposition. a sing methool Solution Ž Note a In method of - SuperNowy deflection & 4! - F(d+9): 1 Jc - - BEIT : we know beam free and that deffection carrying point with Span of Contilever LDeflection at a written due to UOL i 9 - - [ 4+ Go (4) 3 : we know that Contileves beam Carrying UDL Loard over span d thenTotal de flection at ca from h and = quation y = y + y = -f(A+)3 y W13 (34+74) BEI 24 EI y = - [F (1+q) 3 I JE WL3 (31+ 497 2